A coin purse contains 19 coins in nickels, dimes, and quarters. There are three times as many quarters as dimes. The total value of all the coins is $2.90. How many nickels, dimes, and quarters are in the coin purse?
I've gotten this far:
Let x equal the nickels
Let y equal the dimes
Let Z equal the quarters
X+Y+Z = 19
X+y+3z = 2.90
I can't figure out where to go from there.

Answer by rwm(914)

Let x equal the nickels
Let y equal the dimes
Let Z equal the quarters
X+Y+Z = 19
X+y+3z = 2.90
good start but you went awry.
better for memory to use n for nickels d for dimes and q for quarters
n+d+q=19
5n+10d+25q=290
Remember three unknowns need three equations
We have some more info
q=3d
all answers must be positive whole numbers
d=3 n=7 q=9